# Leveraging Skills from Unlabeled Prior Data for Efficient Online Exploration

### v2

@@ Abstract
Unsupervised pretraining has been transformative in many supervised domains. However, applying such ideas to reinforcement learning (RL) presents a unique challenge in that fine-tuning does not involve mimicking task-specific data, but rather exploring and locating the solution through iterative self-improvement. In this work, we study how unlabeled prior trajectory data can be leveraged to learn efficient exploration strategies. While prior data can be used to pretrain a set of low-level skills, or as additional off-policy data for online RL, it has been unclear how to combine these ideas effectively for online exploration. Our method SUPE (Skills from Unlabeled Prior data for Exploration) demonstrates that a careful combination of these ideas compounds their benefits. Our method first extracts low-level skills using a variational autoencoder (VAE), and then pseudo-relabels unlabeled trajectories using an optimistic reward model, transforming prior data into high-level, task-relevant examples. Finally, SUPE uses these transformed examples as additional off-policy data for online RL to learn a high-level policy that composes pretrained low-level skills to explore efficiently. We empirically show that SUPE reliably outperforms prior strategies, successfully solving a suite of long-horizon, sparse-reward tasks. Code: https://github.com/rail-berkeley/supe.
Figure 1: SUPE utilizes unlabeled trajectory data twice, both for offline unsupervised skill pretraining and for online high-level policy learning using RL. Left: in the offline pretraining phase (Stage 1), we unsupervisedly learn both a trajectory segment encoder (a) and a low-level latent conditioned skill policy (b) via a behavior cloning objective where the policy is optimized to reconstruct the action in the trajectory segment. Right: in the online exploration phase (Stage 2), the pretrained trajectory segment encoder (a) and an optimistic reward module (d) are used to pseudo-label the prior data and transform it into high-level trajectories (f) that can be readily consumed by a high-level off-policy RL agent. Leveraging these offline trajectories and the online replay buffer (e), we learn a high-level policy (c) that picks the pretrained low-level skills online to explore in the environment. Finally, the observed transitions and reward values are used to update the optimistic reward module and the online replay buffer.
@@ 1 Introduction
Unsupervised pretraining has been transformative in many supervised domains, such as language (Devlin et al., 2018) and vision (He et al., 2022). Pretrained models can adapt with small numbers of examples, and with better generality (Radford et al., 2019; Brown et al., 2020). However, in contrast to supervised learning, reinforcement learning (RL) presents a unique challenge in that fine-tuning does not involve further mimicking task-specific data, but rather exploring and locating the solution through iterative self-improvement. Thus, the key challenge to address in pretraining for RL is not simply to learn good representations, but to learn an effective exploration strategy for solving downstream tasks.
@@
Pretraining benefits greatly from the breadth of the data. Unlabeled trajectories (i.e., those collected from previous policies whose objectives are unknown) are the most abundantly available, but using them to solve specific tasks can be difficult. It is not enough to simply copy behaviors, which can differ greatly from the current task. There is an entanglement problem – general knowledge of the environment is mixed in with task-specific behaviors. A concrete example is learning from unlabeled locomotion behavior: we wish to learn how to move around the world, but not necessarily to the locations present in the pretraining data. We will revisit this setting in the experimental section.
@@
The entanglement problem can be alleviated through hierarchical decomposition. Specifically, trajectories can be broken into segments of task-agnostic skills, which are composed in various ways to solve various objectives. We posit that unlabeled trajectories thus present a twofold benefit, (1) as a way to learn a diverse set of skills, and (2) as off-policy examples of composing such skills. Notably, prior online RL methods that leverage pretrained skills largely ignore the second benefit, and discard the prior trajectories after the skills are learned (Ajay et al., 2021; Pertsch et al., 2021; Hu et al., 2023; Chen et al., 2024). We instead argue that such trajectories are critical, and can greatly speed up learning. We make use of a simple strategy of learning an optimistic reward model from online samples, and pseudo-relabeling past trajectories with an optimistic reward estimate. The past trajectories can thus be readily utilized as off-policy data, allowing for quick learning even with a very small number of online interactions.
@@
We formalize these insights as SUPE (Skills from Unlabeled Prior data for Exploration), a recipe for maximally leveraging unlabeled prior data in the context of exploration. The prior data is utilized in two capacities, the offline and online phases. In the offline pretraining phase, we extract short segments of trajectories and use them to learn a set of low-level skills. In the online phase, we learn a high-level exploration policy, and again utilize the prior data by labelling each trajectory segment with an optimistic reward estimate. By “double-dipping” in this way, we can utilize both the low-level and high-level structure of prior trajectories to enable efficient exploration online.
@@
Our main contribution is a simple method that leverages unlabeled prior trajectory data to both pretrain skills offline and compose these skills efficiently online for exploration. We instantiate SUPE with a variational autoencoder (VAE) to extract low-level skills, and an off-the-shelf off-policy RL algorithm (Ball et al., 2023) to learn a high-level policy from both online and offline data (Figure 1). Our empirical evaluations on a set of challenging sparse reward tasks show that leveraging the unlabeled prior data during both offline and online learning is crucial for efficient exploration, enabling SUPE to find the sparse reward signal more quickly and achieve more efficient learning over all prior methods (none of which are able to utilize the data both online and offline).
@@ 2 Related Work
Unsupervised skill discovery. Unsupervised skill discovery methods first began in the online setting, where RL agents were tasked with learning structured behaviors in the absence of reward signal (Gregor et al., 2016; Bacon et al., 2017; Florensa et al., 2017; Achiam et al., 2018; Eysenbach et al., 2018; Sharma et al., 2020; Hansen et al., 2020; Park et al., 2023b). These insights naturally transferred to the offline setting as a method of dealing with unlabeled trajectory data. Offline skill discovery methods largely comprise of two categories, those who extract skills based on optimizing unsupervised reward signals (in either the form of policies (Touati et al., 2022; Hu et al., 2023; Frans et al., 2024; Park et al., 2024b) or Q-functions (Chen et al., 2024)), and those who utilize conditional behavior-cloning over subsets of trajectories (Paraschos et al., 2013; Merel et al., 2018; Shankar & Gupta, 2020; Ajay et al., 2021; Singh et al., 2021; Pertsch et al., 2021; Nasiriany et al., 2022). Closest to our method in implementation are Ajay et al. (2021) and Pertsch et al. (2021), who utilize a trajectory-segment VAE to learn low-level skills, and learn a high-level policy online. In contrast to prior methods which all utilize offline data purely for skill-learning and do not keep it around during online training, we show that utilizing the data via relabeling is critical for fast exploration.
@@
Offline to online reinforcement learning. The offline-to-online reinforcement learning methods (Xie et al., 2021b; Song et al., 2023; Lee et al., 2022; Agarwal et al., 2022; Zhang et al., 2023; Zheng et al., 2023; Ball et al., 2023; Nakamoto et al., 2024) focus on efficient online learning with the presence of offline data (often labeled with the reward value). Many offline RL approaches can be applied to this setting – simply run offline RL first on the offline data to convergence as an initialization and then continue training for online learning (using the combined dataset that consists of both offline and online data) (Kumar et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2024). However, such approaches often result in slow online improvements as offline RL objectives tend to overly constrain the policy behaviors to be close to the prior data, limiting the exploration capability. On the other hand, off-policy online RL methods can also be directly applied in this setting by directly treating the offline data as additional off-policy data in the replay buffer and learning the policy from scratch (Lee et al., 2022; Song et al., 2023; Ball et al., 2023). While related in spirit, these methods cannot be directly used in our setting as they require offline data to have reward labels.
@@
Data-driven exploration. A common approach for online exploration is to augment reward bonuses to the perceived rewards and optimize the RL agent with respect to the augmented rewards (Stadie et al., 2015; Bellemare et al., 2016; Houthooft et al., 2016; Pathak et al., 2017; Tang et al., 2017; Ostrovski et al., 2017; Achiam & Sastry, 2017; Merel et al., 2018; Burda et al., 2018; Ermolov & Sebe, 2020; Guo et al., 2022; Lobel et al., 2023). While most exploration methods operate in the purely online setting and focus on adding bonuses to the online replay buffer, recent works also start to explore a more data-driven approach that makes use of an unlabeled prior data to guide online exploration. Li et al. (2024) explore adding bonuses to the offline data, allowing them to optimize the RL agent to be optimistic about states in the data, encouraging exploration around the offline data distribution. Our method explores a similar idea of adding bonuses to the offline data but for training a high-level policy, allowing us to compose pretrained skills effectively for exploration. Hu et al. (2023) explore a slightly different strategy of learning a number of policies using offline RL that each optimizes for a random reward function. Then, it samples actions from these policies online to form an action pool from which the online agent can choose to select for exploration. This approach does not utilize offline data during the online phase and require all the policies (for every random reward function) to be represented separately. In contrast, our method makes use of the offline data as off-policy data for updating the high-level policy and our skills are represented using a single network (with the skill latent being the input to our network). As we will show in our experiments, being able to use offline data is crucial for learning to explore in the environment efficiently.
@@
Hierarchical reinforcement learning. The ability of RL agents to explore and behave effectively over a long horizon is an important research goal in the field of hierarchical RL (HRL) (Dayan & Hinton, 1992; Dietterich, 2000; Vezhnevets et al., 2016; Daniel et al., 2016a; Kulkarni et al., 2016; Vezhnevets et al., 2017; Peng et al., 2017; Riedmiller et al., 2018; Nachum et al., 2018; Ajay et al., 2021; Shankar & Gupta, 2020; Pertsch et al., 2021; Gehring et al., 2021; Xie et al., 2021a). HRL methods typically learn a high-level policy to leverage a space of low-level primitive policies online. These primitives can be either manually specified (Dalal et al., 2021) or pre-trained using unsupervised skill discovery methods as discussed above. While many existing works learn or fine-tune the primitives online along with the high-level policy (Dietterich, 2000; Kulkarni et al., 2016; Vezhnevets et al., 2016; 2017; Nachum et al., 2018; Shankar & Gupta, 2020), others opt for a less flexible but simpler formulation where the primitives are kept fixed after an initial pre-training phase and only the high-level policy is being learned online (Peng et al., 2017; Riedmiller et al., 2018; Ajay et al., 2021; Pertsch et al., 2021; Gehring et al., 2021). Our work adopts the later strategy where we offline pre-train skills using a static, unlabeled dataset. None of prior HRL methods simultaneously leverage offline data for skill pre-training and as additional off-policy data for high-level policy learning online. As we show in our experiments, both of them are crucial in achieving sample efficient learning on challenging sparse-reward tasks.
Options framework. Many existing works on building hierarchical agents also adopt the options framework (Sutton et al., 1999; Menache et al., 2002; Chentanez et al., 2004; Mannor et al., 2004; Şimşek & Barto, 2004; 2007; Konidaris, 2011; Daniel et al., 2016a; Srinivas et al., 2016; Daniel et al., 2016b; Fox et al., 2017; Bacon et al., 2017; Kim et al., 2019; Bagaria & Konidaris, 2019; Bagaria et al., 2024). Different from the approach we take that learns latent skills with a fixed time horizon ($H=4$ in all our experiments), the options framework provides a more flexible way to learn skills with varying time horizon, often defined by learnable initiation and/or termination conditions (Sutton et al., 1999). We opt for the simplified skill definition because it allows us to bypass the need to learn initiation or termination conditions, and frame the skill pretraining phase as a simple supervised learning task.
@@ 3 Problem Formulation
We consider a Markov decision process (MDP) =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathbf{P},\gamma,r,\rho\}$ where $\mathcal{S}$ is the set of all possible states, $\mathcal{A}$ is the set of all possible actions that a policy $\pi(a|s):\mathcal{S}\mapsto\mathcal{P}(\mathcal{A})$ may take, $\mathbf{P}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathcal{P}(\mathcal{S})$ is the transition function that describes the probability distribution over the next state $s^{\prime}$ given the current state and the action taken at the state, $\gamma$ is the discount factor, $r(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, and $\rho:\mathcal{P}(\mathcal{S})$ is the initial state distribution. We have access to a dataset of transitions that are collected from the same MDP with no reward labels: $\mathcal{D}=\{(s_{i},a_{i},s^{\prime}_{i})\}$. During online learning, the agent may interact with the environment by taking actions and observes the next state and the reward specified by transition function $\mathcal{P}$ and the reward function $r$. We aim to develop a method that can leverage the dataset $\mathcal{D}$ to efficiently explore in the MDP to collect reward information, and outputs a well-performing policy $\pi(a|s)$ that achieves good cumulative return in the environment $\eta(\pi)=\mathbb{E}_{\{s_{0}\sim\rho,a_{t}\sim\pi(a_{t}|s_{t}),s_{t+1}\sim\mathbf{P}(\cdot|s_{t},a_{t})\}}\sum_{t=0}^{\infty}\left[\gamma^{t}r(s_{t},a_{t})\right]$. Note that this is different from the zero-shot RL setting (Touati et al., 2022) where the reward function is specified for the online evaluation (only unknown during the unspervised pretraining phase). In our setting, the agent has zero knowledge of the reward function and must actively explore in the environment to identify the task it needs to solve by receiving the reward through environment interactions.
4 Skills from Unlabeled Prior data for Exploration (SUPE)
In this section, we describe in detail how we utilize the unlabeled trajectory dataset to accelerate online exploration. Our method, SUPE, can be roughly divided into two parts. The first part is an offline pretraining phase where we extract skills from the unlabeled prior data with an trajectory-segment VAE. The second part is the online learning phase where we train a high-level off-policy agent to compose the pretrained skills leveraging examples from both prior data and online replay buffer. Algorithm 1 describes our method.

### v4

@@ Abstract
Unsupervised pretraining has been transformative in many supervised domains. However, applying such ideas to reinforcement learning (RL) presents a unique challenge in that fine-tuning does not involve mimicking task-specific data, but rather exploring and locating the solution through iterative self-improvement. In this work, we study how unlabeled offline trajectory data can be leveraged to learn efficient exploration strategies. While prior data can be used to pretrain a set of low-level skills, or as additional off-policy data for online RL, it has been unclear how to combine these ideas effectively for online exploration. Our method SUPE (Skills from Unlabeled Prior data for Exploration) demonstrates that a careful combination of these ideas compounds their benefits. Our method first extracts low-level skills using a variational autoencoder (VAE), and then pseudo-labels unlabeled trajectories with optimistic rewards and high-level action labels, transforming prior data into high-level, task-relevant examples that encourage novelty-seeking behavior. Finally, SUPE uses these transformed examples as additional off-policy data for online RL to learn a high-level policy that composes pretrained low-level skills to explore efficiently. In our experiments, SUPE consistently outperforms prior strategies across a suite of 42 long-horizon, sparse-reward tasks. Code: https://github.com/rail-berkeley/supe.
Machine Learning, ICML
Figure 1: Left: SUPE utilizes unlabeled offline data twice, both for offline unsupervised skill pretraining and online RL for learning the high-level policy. We pseudo-label the unlabeled data with upper-confidence bound (UCB) estimates of the true rewards (reward label: $\hat{r}$) and high-level action estimates (high-level action label: $\hat{z}$). This transforms the unlabeled data into high-level labeled off-policy data that encourages exploration, allowing significantly more sample efficient learning online. Right: Empirically, SUPE attains strong performance across 42 challenging sparse reward tasks (see details in Section 5).
@@ 1 Introduction
Unsupervised pretraining has been transformative in many supervised domains, such as language (Devlin et al., 2018) and vision (He et al., 2022). Pretrained models can be adapted efficiently to new tasks with few examples, and often generalize better than models trained from scratch (Radford et al., 2019; Brown et al., 2020). However, in contrast to supervised learning, reinforcement learning (RL) presents a unique challenge in that fine-tuning does not involve further mimicking task-specific data, but rather exploring and finding the solution through iterative self-improvement. Thus, the key challenge to address in pretraining for RL is not simply to learn good representations, but to learn an effective exploration strategy for solving downstream tasks.
@@
Pretraining benefits greatly from the breadth of the data. Unlabeled trajectories (i.e., those collected from previous policies whose objectives are unknown, or from task-agnostic data collection policies) are the most abundantly available, but using them to solve specific tasks can be difficult. It is not enough to simply copy behaviors, which can differ greatly from the current task. There is an entanglement problem – general knowledge of the environment is mixed in with task-specific behaviors. A concrete example is learning from unlabeled locomotion behavior: we wish to learn how to move around the world, but not necessarily to the locations present in the pretraining data. We will revisit this setting in the experimental section.
@@
The entanglement problem can be alleviated through hierarchical decomposition. Specifically, trajectories can be broken into segments of task-agnostic skills, which are composed in various ways to solve various objectives. We posit that unlabeled trajectories thus present a twofold benefit, (1) as a way to learn a diverse set of skills, and (2) as off-policy examples of composing such skills. Notably, prior online RL methods that leverage pretrained skills ignore the second benefit, and discard the offline trajectories after the skills are learned (Ajay et al., 2021; Pertsch et al., 2021; Hu et al., 2023; Chen et al., 2024). We instead argue that such trajectories are critical, and can greatly speed up learning. We make use of a simple strategy of learning an optimistic reward model from online samples, and pseudo-labeling past trajectories with an optimistic reward estimate. The past trajectories can thus be readily utilized as off-policy data, allowing for quick learning even with a very small number of online interactions.
@@
We formalize these insights as SUPE (Skills from Unlabeled Prior data for Exploration), a recipe for maximally leveraging unlabeled offline data for online exploration (Figure 1). The offline data is utilized during both the offline and online phases. 1 In the offline phase, we extract short segments of trajectories from the unlabeled offline data and use them to pretrain a space of low-level skills represented by a latent conditioned policy $\pi_{\theta}(a|s,z)$. 2 In the online phase, we utilize the unlabeled offline data as additional off-policy data for learning a high-level exploration policy $\pi_{\psi}(z|s)$ by pseudo-labeling each trajectory segment with an upper-confidence bound (UCB) reward estimate of the true rewards and a high-level action estimate ($\hat{z}$). By using offline data in these two ways, we utilize both its low and high-level structure to accelerate online exploration.
@@
Our main contribution is a simple method that leverages unlabeled offline trajectory data to both pretrain skills offline and compose these skills for efficient online exploration. We instantiate SUPE with a variational autoencoder (VAE) to extract low-level skills, and an off-the-shelf off-policy RL algorithm (Ball et al., 2023) to learn a high-level policy from both online and offline data. Our empirical evaluations on a set of challenging sparse reward tasks show that leveraging the unlabeled offline data during both offline and online learning is crucial for efficient exploration, enabling SUPE to find sparse reward signals more quickly and achieve more efficient learning compared to all prior methods (none of which utilize the offline data during both online and offline learning).
@@ 2 Related Work
Unsupervised skill discovery methods were first investigated in the online setting, where RL agents were tasked with learning structured behaviors in the absence of a reward signal (Gregor et al., 2016; Bacon et al., 2017; Florensa et al., 2017; Achiam et al., 2018; Eysenbach et al., 2018; Sharma et al., 2020; Hansen et al., 2020; Campos et al., 2020; Laskin et al., 2021, 2022; Park et al., 2023b; Yang et al., 2023; Bai et al., 2024). These insights naturally transferred to the offline setting as a method of pretraining on unlabeled trajectory data. Offline skill discovery methods largely comprise of two categories, those who extract skills based on optimizing unsupervised reward signals (in either the form of policies (Touati et al., 2022; Hu et al., 2023; Frans et al., 2024; Park et al., 2024b) or Q-functions (Chen et al., 2024)), and those who utilize conditional behavior-cloning over subsets of trajectories (Paraschos et al., 2013; Merel et al., 2018; Shankar & Gupta, 2020; Ajay et al., 2021; Singh et al., 2021; Pertsch et al., 2021; Nasiriany et al., 2022). Closest to our method are Ajay et al. (2021) and Pertsch et al. (2021), who utilize a trajectory-segment VAE to pretrain low-level skills offline and learn a high-level policy online to compose them. In contrast to these methods that utilize offline data purely for offline skill-learning, our method also utilizes the data during online learning via pseudo-labeling. As we show in our experiments, utilizing both the low-level and high-level structure of the offline trajectories is critical for fast exploration and learning online.
@@
Offline-to-online reinforcement learning methods (Xie et al., 2021b; Song et al., 2023; Lee et al., 2022; Agarwal et al., 2022; Zhang et al., 2023; Zheng et al., 2023; Ball et al., 2023; Nakamoto et al., 2024) focus on using task-specific offline data to accelerate online learning. Many offline RL approaches can be applied to this setting — simply run offline RL first on the offline data to convergence as an initialization and then continue training for online learning (using the combined dataset that consists of both offline and online data) (Kumar et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2024). Such approaches often result in slow online improvements as offline RL objectives tend to overly constrain the policy behaviors to be close to the offline data, impairing exploration. On the other hand, off-policy online RL methods can also be directly applied in this setting by treating the offline data as additional off-policy data in the replay buffer and learning the policy from scratch (Lee et al., 2022; Song et al., 2023; Ball et al., 2023). While related in spirit, these methods cannot be directly used in our setting as they require offline data to have reward labels.
@@
Data-driven exploration. A common approach for online exploration is to augment rewards with a bonus that incentives exploration and optimize the agent with respect to these modified rewards (Stadie et al., 2015; Bellemare et al., 2016; Houthooft et al., 2016; Pathak et al., 2017; Tang et al., 2017; Ostrovski et al., 2017; Achiam & Sastry, 2017; Merel et al., 2018; Burda et al., 2018; Ermolov & Sebe, 2020; Guo et al., 2022; Lobel et al., 2023). While most exploration methods operate in the purely online setting and focus on adding bonuses to the online replay buffer, recent works have begun to explore a more data-driven approach that makes use of unlabeled offline data to guide online exploration. Li et al. (2024) added bonuses to the offline data, incentivizing exploration around the offline data distribution. While prior methods apply the bonus directly to dataset transitions, we instead add bonuses to high-level transitions, allowing us to learn a policy that composes pretrained skills effectively for exploration. Hu et al. (2023) proposes a different strategy, first learning a number of policies using offline RL that each optimizes for a random reward function, then sampling from these policies online to achieve structured exploratory behavior. This approach does not utilize offline data during the online phase and requires each policy corresponding to a different reward function to be represented separately. In contrast, our method makes use of the offline data as off-policy data for updating the high-level policy and our skills are represented using a single network parameterized by a skill latent. As we will show, using offline data online is crucial for efficient exploration.
@@
Hierarchical reinforcement learning (HRL). The ability of RL agents to solve long-horizon exploration tasks is an important research goal in the field of HRL (Dayan & Hinton, 1992; Dietterich, 2000; Vezhnevets et al., 2016; Daniel et al., 2016; Kulkarni et al., 2016; Vezhnevets et al., 2017; Peng et al., 2017; Riedmiller et al., 2018; Nachum et al., 2018; Ajay et al., 2021; Shankar & Gupta, 2020; Pertsch et al., 2021; Gehring et al., 2021; Xie et al., 2021a). HRL methods typically learn a high-level policy to leverage a space of low-level primitive policies online. These primitives can be either manually specified (Dalal et al., 2021), learned online (Dietterich, 2000; Kulkarni et al., 2016; Vezhnevets et al., 2016, 2017; Nachum et al., 2018), or pretrained using unsupervised skill discovery methods as discussed above. While many existing works learn or fine-tune the primitives online along with the high-level policy (Dietterich, 2000; Kulkarni et al., 2016; Vezhnevets et al., 2016, 2017; Nachum et al., 2018; Shankar & Gupta, 2020), others opt for a simpler formulation where the primitives are fixed after an initial pre-training phase and only the high-level policy is learned online (Peng et al., 2017; Riedmiller et al., 2018; Ajay et al., 2021; Pertsch et al., 2021; Gehring et al., 2021). Our work adopts the later strategy where we pre-train skills offline using a static, unlabeled dataset. No prior HRL method simultaneously leverages offline data for skill pre-training and as additional off-policy data for learning a high-level policy online. As we show in our experiments, using these approaches in combination is crucial to achieving sample efficient learning on challenging sparse-reward tasks.
See more related work discussions on options and unsupervised pre-training in Appendix D.
@@ 3 Problem Formulation
In this paper, we consider a Markov decision process (MDP) =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathbf{P},\gamma,r,\rho\}$ where $\mathcal{S}$ is the set of all possible states, $\mathcal{A}$ is the set of all possible actions that a policy $\pi(a|s):\mathcal{S}\mapsto\mathcal{P}(\mathcal{A})$ may take, $\mathbf{P}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathcal{P}(\mathcal{S})$ is the transition function that describes the probability distribution over the next state $s^{\prime}$ given the current state and the action taken at the state, $\gamma$ is the discount factor, $r(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, and $\rho:\mathcal{P}(\mathcal{S})$ is the initial state distribution. We have access to a dataset of trajectories that are collected from the same MDP with no reward labels: $\mathcal{D}:=\{\tau^{1},\tau^{2},\cdots\}$ (where $\tau^{i}:=\{s^{i}_{0},a^{i}_{0},s^{i}_{1},a^{i}_{1},s^{i}_{2},\cdots\}$). During online learning, the agent may interact with the environment by taking actions and observes the next state and the reward specified by transition function $\mathcal{P}$ and the reward function $r$. We aim to develop a method that can leverage the dataset $\mathcal{D}$ to efficiently explore in the MDP to collect reward information, and outputs a well-performing policy $\pi(a|s)$ that achieves high cumulative return in the environment $\eta(\pi):=\mathbb{E}_{\{s_{0}\sim\rho,a_{t}\sim\pi(a_{t}|s_{t}),s_{t+1}\sim\mathbf{P}(\cdot|s_{t},a_{t})\}}\sum_{t=0}^{\infty}\left[\gamma^{t}r(s_{t},a_{t})\right]$. Note that this is different from the zero-shot RL setting (Touati et al., 2022) where the reward function is specified for the online evaluation (only unknown during the unsupervised pretraining phase). In our setting, the agent has no knowledge of the reward function a priori and must actively explore the environment online to receive reward information. In addition, our setting is also different from an adjacent setting where offline data only contain observations (Ma et al., 2022; Ghosh et al., 2023; Song et al., 2024) (see Appendix D for more discussions).